OpenStudy (anonymous):

plz check this What is the simplified form of 3/x+2/y divided by 3/x-2/y i got 2y-3x/2y+3x

OpenStudy (anonymous):

@austinL @Paynesdad

OpenStudy (anonymous):

So you have \[\frac{\frac{3}{x} + \frac{2}{y}}{\frac{3}{x} - \frac{2}{y}} = \frac{2y - 3x}{2y + 3x}\]

OpenStudy (anonymous):

yes the first part is the question and the for my answer

OpenStudy (anonymous):

Yes, I think you are very close, but you may have switched numbers somewhere. Also, it might be simplified more but I have to look into it :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

If you multiply nominator and denominator by x, you get?

OpenStudy (anonymous):

i get 3 becuase the x cancels out

OpenStudy (anonymous):

That is a logical first step, as fractions in fractions look ugly. Correct, but remember, to keep the fraction the same you need to multiply every factor with x: \[x\left(\frac{\frac{3}{x} + \frac{2}{y}}{\frac{3}{x} - \frac{2}{y}}\right)\]

OpenStudy (anonymous):

So top and bottom:\[\frac{x\left(\frac{3}{x} + \frac{2}{y}\right)}{x\left(\frac{3}{x} - \frac{2}{y}\right)}\]

OpenStudy (anonymous):

Try the equation editor to make it nicely readable here :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

will you be here in about an hour because i need help but i have to go do something that i wont be back nfor an hour

OpenStudy (anonymous):

I should be yes :)